# Optimal Neural Compressors for the Rate-Distortion-Perception Tradeoff

### v1

@@ Appendix C Proofs
$\leq\frac{\epsilon}{8\beta}+\frac{1}{\sqrt{n}}\sqrt{2\frac{(\sigma^{2}-\nu)\nu}{\sigma^{2}}D_{\mathsf{KL}}(\bm{u}||\tilde{\bm{z}})}$ (83)
$\leq\frac{\epsilon}{8\beta}+\frac{\epsilon}{8\beta}$ (84)
$=\frac{\epsilon}{4\beta}$ (85)
@@
where (82) holds by Ling and Belfiore [2014, Lemma 9] for $n$ sufficiently large, and data processing inequality for $W_{2}$ [Santambrogio, 2015, Lemma 5.2], (83) holds by Talagrand [1996], and (84) holds since
@@
where we use the fact that $\frac{1}{n}\operatorname{\mathbb{E}}[\|\bm{u}\|^{2}]=G(\Lambda^{(n)})\cdot 2\pi e\cdot\frac{(\sigma^{2}-\nu)\nu}{\sigma^{2}}(1+\epsilon_{2})$ and the lattice sequence is sphere-bound-achieving.
For term $B$, let us first divide by $\beta$ and analyze the squared 2-Wasserstein; this gives us $\frac{1}{n}W_{2}^{2}(P_{Q(\alpha\tilde{\bm{x}})+\bm{u}},P_{\bm{y}+\bm{u}})\leq\frac{1}{n}W_{2}^{2}(P_{Q(\alpha\tilde{\bm{x}})},P_{\bm{y}})$ by Santambrogio [2015, Lemma 5.2]. Let $\pi$ be the coupling between $P_{Q(\alpha\tilde{\bm{x}})},P_{\bm{y}}$ induced by the joint $P_{\tilde{\bm{x}},\bm{y}}$; i.e., $\hat{\bm{y}},\bm{y}\sim\pi$ means that $\hat{\bm{y}}=Q(\alpha(\bm{y}+\bm{z}))$ with $\bm{z}\sim\mathcal{N}(0,\nu I_{n})$ as defined above. Then,
@@
where we use the fact that $y,z$ are independent. Note that $\frac{1}{n}\operatorname{\mathbb{E}}[\|\bm{y}\|^{4}]\leq 3(\sigma^{2}-\nu)^{2}$ for $n$ sufficiently large [Zhao and Qian, 2024, Micciancio and Regev, 2004], $\operatorname{\mathbb{E}}[\|\bm{z}\|^{4}]=\nu^{2}n(n+2)$, and $\operatorname{\mathbb{E}}[\langle\bm{y},\bm{z}\rangle^{2}]\leq\operatorname{\mathbb{E}}[\|\bm{y}\|^{2}\|\bm{z}\|^{2}]\leq\operatorname{\mathbb{E}}[\|\bm{y}\|^{2}]\operatorname{\mathbb{E}}[\|\bm{z}\|^{2}]$ by Cauchy-Schwarz and independence. Therefore
@@
$\frac{1}{n}\sqrt{\operatorname{\mathbb{E}}_{\bm{y},\bm{z}}\left[\|\hat{\bm{y}}-\bm{y}\|^{4}\right]\operatorname{\mathbb{P}}(\mathcal{E})}\leq\frac{\epsilon^{2}}{16\beta^{2}},$ (105)
for $n$ sufficiently large, since the error of the MAP estimate satisfies $\lim_{n\rightarrow\infty}\operatorname{\mathbb{P}}(\mathcal{E})=0$ (exponentially fast) [Ling and Belfiore, 2014, Lemma 11] for AWGN-good lattices with the volume $V(\Lambda^{(n)})$ we chose above. Hence
$\frac{1}{\sqrt{n}}W_{2}(P_{\beta(Q(\alpha\tilde{\bm{x}})+\bm{u})},P_{\beta(\bm{y}+\bm{u})})$ $\leq\frac{\beta}{\sqrt{n}}W_{2}(P_{Q(\alpha\tilde{\bm{x}})},P_{\bm{y}})$ (106)
@@
(105)${{\leq}}\beta\frac{\epsilon}{4\beta}=\frac{\epsilon}{4}.$ (108)
We note here that [Ling and Belfiore, 2014, Lemma 11] has a signal-to-noise ratio (SNR) condition of $\frac{\sigma^{2}-\nu}{\nu}>e$ for the error probability to decay exponentially fast. This was removed in Campello et al. [2018] via deterministic dithering; there exists $\bm{t}\in\mathbb{R}^{n}$ with $g_{a}^{(n)}(\bm{v})=\alpha\bm{v}-\bm{t}$, and $g_{s}^{(n)}(\bm{v})=\beta(\bm{v}+\bm{t})$ such that the error probability results hold with no SNR condition. While we do not explicitly use these arguments in the remainder of the proof, they can easily be extended via Campello et al. [2018, Thm. 1].
Next, we address the distortion term. Fix $\epsilon>0$. We have that
@@
$\qquad+\underbrace{\frac{1}{n}\operatorname{\mathbb{E}}_{\tilde{\bm{x}},\bm{u}}\left[\|\tilde{\bm{x}}-\beta(Q(\alpha\tilde{\bm{x}})-\bm{u})\|^{2}\right]}_{S_{2}}.$ (109)
By Ling and Belfiore [2014], Regev [2009], for any $\epsilon^{\prime\prime}>0$, $|dP_{\tilde{\bm{x}}}(\bm{x})-dP_{X^{n}}(\bm{x})|\leq\epsilon^{\prime\prime}dP_{X^{n}}(\bm{x}),\forall\bm{x}$, by the choice of the lattice sequence $\Lambda^{(n)}$, for $n$ sufficiently large. The first term $S_{1}$ can be written as
@@
$\frac{1}{n}\operatorname{\mathbb{E}}_{X^{n},\bm{u}}\left[\|X^{n}-\beta(Q(\alpha X^{n})-\bm{u})\|^{2}\right]\leq\frac{1}{1-\epsilon^{\prime\prime}}S_{2}.$ (111)
@@
$S_{2}$ $=\frac{1}{n}\operatorname{\mathbb{E}}\left[\|\tilde{\bm{x}}-\beta(Q(\alpha\tilde{\bm{x}})-\bm{u})\|^{2}\mathbbm{1}\{\mathcal{E}^{\complement}\}\right]+\frac{1}{n}\operatorname{\mathbb{E}}\left[\|\tilde{\bm{x}}-\beta(Q(\alpha\tilde{\bm{x}})-\bm{u})\|^{2}\mathbbm{1}\{\mathcal{E}\}\right]$ (112)
$\leq\frac{1}{n}\operatorname{\mathbb{E}}\left[\|\tilde{\bm{x}}-\beta(Q(\alpha\tilde{\bm{x}})-\bm{u})\|^{2}\mathbbm{1}\{\mathcal{E}^{\complement}\}\right]+\frac{1}{n}\sqrt{\operatorname{\mathbb{E}}\left[\|\tilde{\bm{x}}-\beta(Q(\alpha\tilde{\bm{x}})-\bm{u})\|^{4}\right]\operatorname{\mathbb{P}}(\mathcal{E})},$ (113)
@@
$\operatorname{\mathbb{E}}\left[\|\tilde{\bm{x}}-\beta(Q(\alpha\tilde{\bm{x}})-\bm{u})\|^{4}\right]$ (114)
$\leq 8(1-\beta\alpha)\operatorname{\mathbb{E}}\left[\|\tilde{\bm{x}}\|^{4}\right]+8\beta\operatorname{\mathbb{E}}\left[\|Q(\alpha\tilde{\bm{x}})-\alpha\tilde{\bm{x}}\|^{4}\right]+8\beta\operatorname{\mathbb{E}}\left[\|\bm{u}\|^{4}\right].$ (115)
@@
$8\beta\operatorname{\mathbb{E}}\left[\|\bm{u}\|^{4}\right]\leq 8\beta r^{4}_{\mathrm{cov}}(\Lambda^{(n)})=O(n^{2})$ (116)
by Lemma C.3. For the first term,
@@
for $n$ sufficiently large, by the independence of $\bm{y}$ and $\bm{z}$, and using 4th moment results on lattice Gaussians from Zhao and Qian [2024], Micciancio and Regev [2004]. By the choice of $\alpha$, $Q(\alpha\tilde{\bm{x}})$ computes the MAP estimate of $\bm{y}$ [Ling and Belfiore, 2014, Prop. 3], and thus by Ling and Belfiore [2014, Lemma 9], its error probability satisfies $\lim_{n\rightarrow\infty}\operatorname{\mathbb{P}}(\mathcal{E})=0$, where the convergence is exponentially fast in $n$, since the lattice sequence was assumed AWGN-good. Thus $\frac{1}{n}\sqrt{\operatorname{\mathbb{E}}\left[\|\tilde{\bm{x}}-\beta(Q(\alpha\tilde{\bm{x}})-\bm{u})\|^{4}\right]\operatorname{\mathbb{P}}(\mathcal{E})}$ vanishes when $n\rightarrow\infty$, and we have
$\lim_{n\rightarrow\infty}S_{2}$ $\leq\lim_{n\rightarrow\infty}\frac{1}{n}\operatorname{\mathbb{E}}\left[\|\tilde{\bm{x}}-\beta(Q(\alpha\tilde{\bm{x}})-\bm{u})\|^{2}\mathbbm{1}\{\mathcal{E}^{\complement}\}\right]$ (121)
$=\lim_{n\rightarrow\infty}\operatorname{\mathbb{E}}\left[\frac{1}{n}\|\tilde{\bm{x}}-\beta(\bm{y}-\bm{u})\|^{2}\mathbbm{1}\{\mathcal{E}^{\complement}\}\right]$ (122)
$\leq\lim_{n\rightarrow\infty}\operatorname{\mathbb{E}}\left[\frac{1}{n}\|\tilde{\bm{x}}-\beta(\bm{y}-\bm{u})\|^{2}\right]\|\mathbbm{1}\{\mathcal{E}^{\complement}\}\|_{\infty}$ (123)
$=\lim_{n\rightarrow\infty}\frac{1}{n}\operatorname{\mathbb{E}}\left[\|\tilde{\bm{x}}-\beta(\bm{y}-\bm{u})\|^{2}\right]$ (124)
$=\lim_{n\rightarrow\infty}\frac{1}{n}\operatorname{\mathbb{E}}\left[\|(1-\beta)\bm{y}+\bm{z}-\beta\bm{u}\|^{2}\right]$ (125)
$=\lim_{n\rightarrow\infty}(1-\beta)^{2}\frac{1}{n}\operatorname{\mathbb{E}}[\|\bm{y}\|^{2}]+\frac{1}{n}\operatorname{\mathbb{E}}[\|\bm{z}\|^{2}]+\beta^{2}\frac{1}{n}\operatorname{\mathbb{E}}[\|\bm{u}\|^{2}]$ (126)
$\leq\lim_{n\rightarrow\infty}(1-\beta)^{2}(\sigma^{2}-\nu)+\nu+\beta^{2}\cdot G(\Lambda^{(n)})\cdot 2\pi e\cdot\frac{(\sigma^{2}-\nu)\nu}{\sigma^{2}}(1+\epsilon_{2})$ (127)
$=(1-\beta)^{2}(\sigma^{2}-\nu)+\nu+\beta^{2}\frac{(\sigma^{2}-\nu)\nu}{\sigma^{2}}$ (128)
$=D,$ (129)
where (123) is by Hölder’s inequality, and (127) holds by Banaszczyk [1993], and since the lattice second moment satisfies $\frac{1}{n}\operatorname{\mathbb{E}}[\|\bm{u}\|^{2}]=G(\Lambda^{(n)})\cdot V(\Lambda^{(n)})^{2/n}$, and the lattice sequence is sphere-bound-achieving. The result follows by combining with (111) and taking $\epsilon^{\prime\prime}\rightarrow 0$ with $n\rightarrow\infty$.
@@
The first term $R_{1}$ will vanish as $n\rightarrow\infty$, due to the following. By Regev [2009, Claim 3.9], we have that for any $\epsilon^{\prime\prime}>0$, $|dP_{\tilde{\bm{x}}}(\bm{x})-dP_{X^{n}}(\bm{x})|\leq\epsilon^{\prime\prime}dP_{X^{n}}(\bm{x}),\forall\bm{x}$, by the choice of the lattice sequence $\Lambda^{(n)}$, for $n$ sufficiently large (see also Ling and Belfiore [2014, Lemma 9]). Therefore,
@@
$\leq\epsilon^{\prime\prime}\frac{1}{n}\int-\log(dP_{Q(\alpha\bm{x})})dP_{X^{n}}(\bm{x})d\bm{x}$ (133)
$=\epsilon^{\prime\prime}\frac{1}{n}H(Q(\alpha X^{n})),$ (134)
@@
$\frac{1}{n}H(Q(\alpha X^{n}))\leq\frac{1}{1-\epsilon^{\prime\prime}}R_{2},$ (135)
@@
For $R_{2}$, this is the per-dimension entropy of $Q(\alpha\tilde{\bm{x}})$. Let $p(\bm{\lambda}):=\Pr\left(Q(\alpha\tilde{\bm{x}})=\bm{\lambda}\right)$ be the PMF of $Q(\alpha\tilde{\bm{x}})$ supported on $\bm{\lambda}\in\Lambda$, and let $q_{\bm{y}}(\bm{\lambda})$ be the lattice Gaussian PMF of $\bm{y}\sim\mathcal{N}_{\Lambda}(\bm{0},\sigma^{2}-\nu)$ defined in Def. C.2. Then
@@
The second moment of $Q(\alpha\tilde{\bm{x}})$ satisfies
$\leq 8\|\bm{y}-\alpha\tilde{\bm{x}}\|^{4}+8\|\alpha\tilde{\bm{x}}\|^{4}$ (146)
where (145) is by triangle inequality and Cauchy-Schwarz, and (146) holds since $Q$ finds the closest vector to $\alpha\tilde{\bm{x}}$. Following (101), we have that the expectations satisfy
$=O(n^{2}),$ (152)
and
$=O(n^{2}),$ (156)
for $n$ sufficiently large. This implies that
and therefore
$\frac{1}{n}\sqrt{\operatorname{\mathbb{E}}\left[\|Q(\alpha\tilde{\bm{x}})\|^{4}\right]\operatorname{\mathbb{P}}(\mathcal{E})}<2\epsilon^{\prime\prime}(\sigma^{2}-\nu)$ (158)
for $n$ sufficiently large since $\lim_{n\rightarrow\infty}\operatorname{\mathbb{P}}(\mathcal{E})=0$ exponentially fast [Ling and Belfiore, 2014].
@@
$\leq\frac{1}{2(\sigma^{2}-\nu)}\left[\frac{1}{n}\operatorname{\mathbb{E}}{\|\bm{y}\|^{2}}+2\epsilon^{\prime\prime}(\sigma^{2}-\nu)\right]+\frac{1}{n}\log\rho_{\sqrt{\sigma^{2}-\nu}}(\Lambda)$ (161)

### v2

@@ Appendix D Proofs
$\frac{1}{\sqrt{n}}W_{2}\left(P_{\hat{X}^{n}},P_{\beta(Q(\alpha\tilde{\bm{x}})+s\bm{u})}\right)$ $\leq\beta\sqrt{8\epsilon_{z}\left(2\frac{\sigma^{2}-\nu}{1-4\epsilon_{z}}+2s^{2}\frac{(\sigma^{2}-\nu)\nu}{\sigma^{2}}\right)}.$ (110)
For the first term on the right of (93), we again apply triangle inequality:
$\frac{1}{\sqrt{n}}W_{2}(P_{X^{n}},P_{\beta(Q(\alpha\tilde{\bm{x}})+s\bm{u})})\leq\underbrace{\frac{1}{\sqrt{n}}W_{2}(P_{X^{n}},P_{\beta(\bm{y}+s\bm{u})})}_{A}+\underbrace{\frac{1}{\sqrt{n}}W_{2}(P_{\beta(Q(\alpha\tilde{\bm{x}})+s\bm{u})},P_{\beta(\bm{y}+s\bm{u})})}_{B}.$ (111)
For term $A$, let $\tilde{\bm{z}}\sim\mathcal{N}\left(0,\frac{(\sigma^{2}-\nu)\nu}{\sigma^{2}}I_{n}\right)$ and $\bm{r}\sim\mathcal{N}\left(0,\frac{\sigma^{2}}{\beta^{2}}I_{n}\right)$. Then
[…]
$\leq\frac{1}{\sqrt{n}}\sqrt{8\epsilon_{u}\operatorname{\mathbb{E}}_{\bm{r}}[\|\bm{r}\|^{2}]}$ (113)
$=\sqrt{8\epsilon_{u}\frac{\sigma^{2}}{\beta^{2}}},$ (114)
where (112) is by Villani [2016, Thm. 6.15], and (113) holds since $s,\beta$ satisfy (88) and by applying Ling and Belfiore [2014, Lemma 9]. Therefore
$\frac{1}{\beta\sqrt{n}}W_{2}(P_{X^{n}},P_{\beta(\bm{y}+s\bm{u})})$ $=\frac{1}{\sqrt{n}}W_{2}\left(P_{\bm{r}},P_{\bm{y}+s\bm{u}}\right)$ (115)
$\leq\frac{1}{\sqrt{n}}W_{2}\left(P_{\bm{r}},P_{\bm{y}+s\tilde{\bm{z}}}\right)+\frac{1}{\sqrt{n}}W_{2}\left(P_{\bm{y}+s\tilde{\bm{z}}},P_{\bm{y}+s\bm{u}}\right)$ (116)
$\leq\sqrt{8\epsilon_{u}\frac{\sigma^{2}}{\beta^{2}}}+s\frac{1}{\sqrt{n}}W_{2}(P_{\bm{u}},P_{\tilde{\bm{z}}})$ (117)
$\leq\sqrt{8\epsilon_{u}\frac{\sigma^{2}}{\beta^{2}}}+s\frac{1}{\sqrt{n}}\sqrt{2\frac{(\sigma^{2}-\nu)\nu}{\sigma^{2}}D_{\mathsf{KL}}(\bm{u}||\tilde{\bm{z}})}$ (118)
$\leq\sqrt{8\epsilon_{u}\frac{\sigma^{2}}{\beta^{2}}}+\frac{\epsilon}{2\beta},$ (119)
@@
where (117) holds by Ling and Belfiore [2014, Lemma 9] for $n$ sufficiently large, and data processing inequality for $W_{2}$ [Santambrogio, 2015, Lemma 5.2], (118) holds by Talagrand [1996], and (119) holds because
@@
where we use the fact that $\frac{1}{n}\operatorname{\mathbb{E}}[\|\bm{u}\|^{2}]=G(\Lambda^{(n)})\cdot 2\pi e\cdot\frac{(\sigma^{2}-\nu)\nu}{\sigma^{2}}(1+\epsilon_{2})$ and the lattice sequence is sphere-bound-achieving. Thus
$A$ $=\frac{1}{\sqrt{n}}W_{2}(P_{X^{n}},P_{\beta(\bm{y}+\bm{u})})\leq\sqrt{8\sigma^{2}\epsilon_{u}}+\frac{\epsilon}{2}.$ (125)
For term $B$, let us first divide by $\beta$ and analyze the squared 2-Wasserstein; this gives us $\frac{1}{n}W_{2}^{2}(P_{Q(\alpha\tilde{\bm{x}})+s\bm{u}},P_{\bm{y}+s\bm{u}})\leq\frac{1}{n}W_{2}^{2}(P_{Q(\alpha\tilde{\bm{x}})},P_{\bm{y}})$ by Santambrogio [2015, Lemma 5.2]. Let $\pi$ be the coupling between $P_{Q(\alpha\tilde{\bm{x}})},P_{\bm{y}}$ induced by the joint $P_{\tilde{\bm{x}},\bm{y}}$; i.e., $\hat{\bm{y}},\bm{y}\sim\pi$ means that $\hat{\bm{y}}=Q(\alpha(\bm{y}+\bm{z}))$ with $\bm{z}\sim\mathcal{N}(0,\nu I_{n})$ as defined above. Then,
@@
where we use the fact that $y,z$ are independent. Note that $\frac{1}{n}\operatorname{\mathbb{E}}[\|\bm{y}\|^{4}]\leq 3(\sigma^{2}-\nu)^{2}$ for $n$ sufficiently large [Zhao and Qian, 2024, Micciancio and Regev, 2004], $\operatorname{\mathbb{E}}[\|\bm{z}\|^{4}]=\nu^{2}n(n+2)$, and $\operatorname{\mathbb{E}}[\langle\bm{y},\bm{z}\rangle^{2}]\leq\operatorname{\mathbb{E}}[\|\bm{y}\|^{2}\|\bm{z}\|^{2}]\leq\operatorname{\mathbb{E}}[\|\bm{y}\|^{2}]\operatorname{\mathbb{E}}[\|\bm{z}\|^{2}]$ by Cauchy-Schwarz and independence. Additionally, $\operatorname{\mathbb{E}}[\|\bm{y}\|^{2}]\leq n(\sigma^{2}-\nu)$ by Banaszczyk [1993]. Therefore
@@
$\frac{1}{n}\sqrt{\operatorname{\mathbb{E}}_{\bm{y},\bm{z}}\left[\|\hat{\bm{y}}-\bm{y}\|^{4}\right]\operatorname{\mathbb{P}}(\mathcal{E})}\leq\frac{\epsilon^{2}}{4\beta^{2}},$ (140)
for $n$ sufficiently large, since the error of the MAP estimate satisfies $\lim_{n\rightarrow\infty}\operatorname{\mathbb{P}}(\mathcal{E})=0$. Hence
$\frac{1}{\sqrt{n}}W_{2}(P_{\beta(Q(\alpha\tilde{\bm{x}})+s\bm{u})},P_{\beta(\bm{y}+s\bm{u})})$ $\leq\frac{\beta}{\sqrt{n}}W_{2}(P_{Q(\alpha\tilde{\bm{x}})},P_{\bm{y}})$ (141)
@@
(140)${{\leq}}\beta\frac{\epsilon}{2\beta}=\frac{\epsilon}{2}.$ (143)
In conclusion, we have shown
$\frac{1}{\sqrt{n}}W_{2}(P_{X^{n}},P_{\hat{X}^{n}})\leq\sqrt{8\sigma^{2}\epsilon_{u}}+\beta\sqrt{8\epsilon_{z}\left(2\frac{\sigma^{2}-\nu}{1-4\epsilon_{z}}+2s^{2}\frac{(\sigma^{2}-\nu)\nu}{\sigma^{2}}\right)}+\epsilon.$ (144)
Next, we address the distortion term. We have that
@@
$\qquad+\underbrace{\frac{1}{n}\operatorname{\mathbb{E}}_{\tilde{\bm{x}},\bm{u}}\left[\|\tilde{\bm{x}}-\beta(Q(\alpha\tilde{\bm{x}})-s\bm{u})\|^{2}\right]}_{S_{2}}.$ (145)
By Regev [2009, Claim 3.9] and Ling and Belfiore [2014, Lemma 9], $|dP_{\tilde{\bm{x}}}(\bm{x})-dP_{X^{n}}(\bm{x})|\leq 4\epsilon_{z}dP_{X^{n}}(\bm{x}),\forall\bm{x}$. The first term $S_{1}$ can be written as
@@
$\frac{1}{n}\operatorname{\mathbb{E}}_{X^{n},\bm{u}}\left[\|X^{n}-\beta(Q(\alpha X^{n})-s\bm{u})\|^{2}\right]\leq\frac{1}{1-4\epsilon_{z}}S_{2}.$ (147)
@@
$S_{2}$ $=\frac{1}{n}\operatorname{\mathbb{E}}\left[\|\tilde{\bm{x}}-\beta(Q(\alpha\tilde{\bm{x}})-s\bm{u})\|^{2}\mathbbm{1}\{\mathcal{E}^{\complement}\}\right]+\frac{1}{n}\operatorname{\mathbb{E}}\left[\|\tilde{\bm{x}}-\beta(Q(\alpha\tilde{\bm{x}})-s\bm{u})\|^{2}\mathbbm{1}\{\mathcal{E}\}\right]$ (148)
$\leq\frac{1}{n}\operatorname{\mathbb{E}}\left[\|\tilde{\bm{x}}-\beta(Q(\alpha\tilde{\bm{x}})-s\bm{u})\|^{2}\mathbbm{1}\{\mathcal{E}^{\complement}\}\right]+\frac{1}{n}\sqrt{\operatorname{\mathbb{E}}\left[\|\tilde{\bm{x}}-\beta(Q(\alpha\tilde{\bm{x}})-s\bm{u})\|^{4}\right]\operatorname{\mathbb{P}}(\mathcal{E})},$ (149)
@@
$\operatorname{\mathbb{E}}\left[\|\tilde{\bm{x}}-\beta(Q(\alpha\tilde{\bm{x}})-s\bm{u})\|^{4}\right]$ (150)
$\leq 8(1-\beta\alpha)\operatorname{\mathbb{E}}\left[\|\tilde{\bm{x}}\|^{4}\right]+8\beta\operatorname{\mathbb{E}}\left[\|Q(\alpha\tilde{\bm{x}})-\alpha\tilde{\bm{x}}\|^{4}\right]+8\beta\operatorname{\mathbb{E}}\left[\|s\bm{u}\|^{4}\right].$ (151)
@@
$8\beta\operatorname{\mathbb{E}}\left[\|s\bm{u}\|^{4}\right]\leq 8\beta s^{4}r^{4}_{\mathrm{cov}}(\Lambda^{(n)})=O(n^{2})$ (152)
by Lemma D.3. For the first term,
@@
for $n$ sufficiently large, by the independence of $\bm{y}$ and $\bm{z}$, and using 4th moment results on lattice Gaussians from Zhao and Qian [2024], Micciancio and Regev [2004]. Since $\lim_{n\rightarrow\infty}\operatorname{\mathbb{P}}(\mathcal{E})=0$, we have that $\frac{1}{n}\sqrt{\operatorname{\mathbb{E}}\left[\|\tilde{\bm{x}}-\beta(Q(\alpha\tilde{\bm{x}})-s\bm{u})\|^{4}\right]\operatorname{\mathbb{P}}(\mathcal{E})}<\frac{\epsilon}{2}$ for $n$ sufficiently large. Thus
$S_{2}$ $\leq\frac{1}{n}\operatorname{\mathbb{E}}\left[\|\tilde{\bm{x}}-\beta(Q(\alpha\tilde{\bm{x}})-s\bm{u})\|^{2}\mathbbm{1}\{\mathcal{E}^{\complement}\}\right]+\frac{\epsilon}{2}$ (157)
$=\operatorname{\mathbb{E}}\left[\frac{1}{n}\|\tilde{\bm{x}}-\beta(\bm{y}-s\bm{u})\|^{2}\mathbbm{1}\{\mathcal{E}^{\complement}\}\right]+\frac{\epsilon}{2}$ (158)
$\leq\operatorname{\mathbb{E}}\left[\frac{1}{n}\|\tilde{\bm{x}}-\beta(\bm{y}-s\bm{u})\|^{2}\right]\|\mathbbm{1}\{\mathcal{E}^{\complement}\}\|_{\infty}+\frac{\epsilon}{2}$ (159)
$=\frac{1}{n}\operatorname{\mathbb{E}}\left[\|\tilde{\bm{x}}-\beta(\bm{y}-s\bm{u})\|^{2}\right]+\frac{\epsilon}{2}$ (160)
$=\frac{1}{n}\operatorname{\mathbb{E}}\left[\|(1-\beta)\bm{y}+\bm{z}-\beta s\bm{u}\|^{2}\right]+\frac{\epsilon}{2}$ (161)
$=(1-\beta)^{2}\frac{1}{n}\operatorname{\mathbb{E}}[\|\bm{y}\|^{2}]+\frac{1}{n}\operatorname{\mathbb{E}}[\|\bm{z}\|^{2}]+s^{2}\beta^{2}\frac{1}{n}\operatorname{\mathbb{E}}[\|\bm{u}\|^{2}]+\frac{\epsilon}{2}$ (162)
$\leq(1-\beta)^{2}(\sigma^{2}-\nu)+\nu+s^{2}\beta^{2}\cdot G(\Lambda^{(n)})\cdot 2\pi e\cdot\frac{(\sigma^{2}-\nu)\nu}{\sigma^{2}}(1+\epsilon_{2})+\frac{\epsilon}{2}$ (163)
$=(1-\beta)^{2}(\sigma^{2}-\nu)+\nu+s^{2}\beta^{2}\frac{(\sigma^{2}-\nu)\nu}{\sigma^{2}}+\epsilon,$ (164)
for $n$ sufficiently large, where (159) is by Hölder’s inequality, and (163) holds by Banaszczyk [1993], and since the lattice second moment satisfies $\frac{1}{n}\operatorname{\mathbb{E}}[\|\bm{u}\|^{2}]=G(\Lambda^{(n)})\cdot V(\Lambda^{(n)})^{2/n}$, and the lattice sequence is quantization-good (i.e., sphere-bound-achieving). The result follows by combining (147) and (164).
@@
The first term $R_{1}$ will vanish as $n\rightarrow\infty$, due to the following. By Regev [2009, Claim 3.9] and Ling and Belfiore [2014, Lemma 9], we have that $|dP_{\tilde{\bm{x}}}(\bm{x})-dP_{X^{n}}(\bm{x})|\leq 4\epsilon_{z}dP_{X^{n}}(\bm{x}),\forall\bm{x}$, by the choice of the lattice sequence $\Lambda^{(n)}$, for $n$ sufficiently large. Therefore,
@@
$\leq 4\epsilon_{z}\frac{1}{n}\int-\log(dP_{Q(\alpha\bm{x})})dP_{X^{n}}(\bm{x})d\bm{x}$ (168)
$=4\epsilon_{z}\frac{1}{n}H(Q(\alpha X^{n})),$ (169)
@@
$\frac{1}{n}H(Q(\alpha X^{n}))\leq\frac{1}{1-4\epsilon_{z}}R_{2},$ (170)
@@
For $R_{2}$, this is the per-dimension entropy of $Q(\alpha\tilde{\bm{x}})$. Let $p(\bm{\lambda}):=\Pr\left(Q(\alpha\tilde{\bm{x}})=\bm{\lambda}\right)$ be the PMF of $Q(\alpha\tilde{\bm{x}})$ supported on $\bm{\lambda}\in\Lambda$, and let $q_{\bm{y}}(\bm{\lambda})$ be the lattice Gaussian PMF of $\bm{y}\sim\mathcal{N}_{\Lambda}(\bm{0},\sigma^{2}-\nu)$. Then
@@
$\leq\frac{1}{2(\sigma^{2}-\nu)}\left[\frac{1}{n}\operatorname{\mathbb{E}}{\|\bm{y}\|^{2}}+2\epsilon(\sigma^{2}-\nu)\right]+\frac{1}{n}\log\rho_{\sqrt{\sigma^{2}-\nu}}(\Lambda)$ (177)